DIFFERENTIAL OPTICAL ABSORPTION SPECTROMETER FOR MONITORING URBAN ATMO-SPHERIC POLLUTANTS AND THEIR DISTRIBUTION

\underline{CHRISTOPHER P. BEEKMAN} and HEATHER C. ALLEN, Environmental Science Graduate Program, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

A differential absorption spectrometer (DOAS) has been developed for monitoring the concentration of urban atmospheric pollutants such as NO_{2}, formaldehyde, and O_{3}. In addition to concentrations and distributions of trace atmospheric species, the DOAS provides information on elastic and inelastic scattering in the atmosphere. Current work on concentration retrieval algorithms and the development of statistical analysis methods will be presented.